# MSDet: Receptive Field Enhanced Multiscale Detection for Tiny Pulmonary Nodule

### v1

@@ IV-E1 Clinical Impact
MSDet’s multiscale detection and attention mechanisms, including ERD and PCAM, address key challenges in identifying small and occluded pulmonary nodules, which are often missed due to their size and complex surroundings. By enhancing detection accuracy and reducing false positives, MSDet assists radiologists in early malignancy identification, boosting diagnostic confidence and enabling timely interventions. Its integration into clinical workflows can streamline lung cancer screening by automating nodule detection, reducing the radiologists’ workload, and allowing them to focus on critical cases. MSDet’s lightweight architecture supports real-time deployment without compromising speed or accuracy, making it well-suited for high-volume imaging environments. As screening programs expand, MSDet can help prioritize urgent cases and ease the burden on healthcare systems.
@@ IV-E2 Contributions to Early Diagnosis and Broader Societal Impact
MSDet’s precise detection of small pulmonary nodules facilitates earlier lung cancer diagnoses, improving survival rates and enabling less invasive treatments. By reducing false positives, it minimizes unnecessary follow-ups, alleviating patient stress and lowering healthcare costs. This enhanced accuracy streamlines the diagnostic process, making healthcare more efficient. Globally, MSDet can have a transformative impact, especially in regions with limited access to radiologists. By providing high-quality automated screening, it helps bridge healthcare disparities, ensuring timely and accurate diagnoses for all patients. Ultimately, MSDet contributes to improved patient outcomes, reduced lung cancer mortality, and lower healthcare burdens.
@@ IV-F Visualization
In this section, we present a visual comparison of the detection results to validate the effectiveness of the proposed MSDet model for lung nodule detection. As illustrated in Figure 10, MSDet consistently outperforms YOLOv4, Scaled_YOLOv4, and the Basic Model across various scenarios, including both single and multiple nodule detections. For instance, MSDet achieved 94% accuracy in detecting nodules in complex imaging backgrounds, while YOLOv4 and Scaled_YOLOv4 exhibited lower accuracies due to false positives and false negatives. In multi-nodule cases, MSDet demonstrated superior precision, accurately identifying all nodules without false positives or false negatives, outperforming the other models by 6% to 8%. These results emphasize the robustness and accuracy of MSDet, especially in handling complex imaging conditions and challenging nodules, making it highly reliable for clinical applications.
Lung parenchyma segmentation further enhances detection accuracy by providing cleaner image backgrounds. This process involves renormalizing the images, using K-means clustering to separate lung tissue from the background, and refining the lung mask through erosion and dilation. As shown in Figure 11, models such as YOLOv4, Scaled_YOLOv4, the Basic Model, and MSDet demonstrate improved performance post-segmentation compared to raw data. MSDet, in particular, achieved up to 98% accuracy in detecting multiple nodules, reducing both false positives and false negatives. Overall, lung parenchyma segmentation significantly boosts detection precision across all models, highlighting its importance in complex imaging tasks.
Figure 11: Visual comparison of lung nodule detection results on the LUNA16 dataset, illustrating the impact of lung parenchyma segmentation on improving detection accuracy. MSDet shows a notable reduction in false positives and an accuracy of up to 98% in complex scenarios, highlighting the effectiveness of segmentation in enhancing the clarity and precision of nodule detection across different algorithms.
@@ V Conclusion
This paper proposes a novel method, MSDet, for lung nodule detection in CT images, specifically designed to address multi-scale nodule detection and ensure accurate identification. To overcome the limitations of traditional algorithms in detecting small nodules, we introduced TODB, which enhances feature extraction for tiny lung nodules. Additionally, we designed an ERD strategy to refine occluded nodule features by constructing an effective receptive field, enabling accurate detection of occluded nodules. Furthermore, we proposed a new attention mechanism, PCAM, which adjusts attention weights based on the importance of different positions and channels. This optimizes feature representation learning and improves adaptability to nodules of varying scales, thereby reducing missed and false detections. In summary, MSDet presents a novel framework for multi-scale tiny nodule detection. Empirical validation demonstrates that our model not only detects small nodules effectively in complex backgrounds but also surpasses existing benchmarks on the public LUNA16 dataset. This study marks a significant advancement in medical image detection, with the potential to profoundly influence future medical image processing endeavors.
Acknowledgment
This work was supported by the National Natural Science Foundation of China (Grant No. 62306027) and Southwest Minzu University. We would like to express our gratitude to our collaborators for their valuable contributions and support throughout this study.

### v2

@@ IV-E1 Clinical Impact
MSDet improves accuracy and reduces false positives, enhancing early malignancy detection. Its automated, lightweight design eases radiologists’ workload and supports real-time deployment, making it ideal for large-scale screenings and alleviating healthcare system burdens.
@@ IV-E2 Contributions to Early Diagnosis and Broader Societal Impact
MSDet enables earlier lung cancer detection, improving survival rates and reducing invasive treatments. By minimizing false positives, it cuts follow-up costs and stress, while enhancing diagnostic efficiency, especially in underserved areas, ultimately reducing healthcare burdens.
@@ IV-F Visualization
We validate the effectiveness of MSDet for lung nodule detection through visual comparison. As shown in Figure 9, MSDet outperforms YOLOv4, Scaled_YOLOv4, and the Basic Model in both single and multi-nodule tasks, achieving 94% accuracy in complex backgrounds. It identifies all nodules with no false positives or negatives, outperforming other models by 6-8%. Lung parenchyma segmentation further boosts accuracy, as shown in Figure 10, enabling MSDet to achieve up to 98% accuracy and significantly reduce false detections. These results demonstrate the robustness of MSDet and the importance of segmentation for improving performance in complex imaging conditions, making it well-suited for clinical applications.
Figure 10: Visual comparison of lung nodule detection on the LUNA16 dataset, showing the impact of lung parenchyma segmentation. MSDet achieves up to 98% accuracy, reducing false positives and improving detection precision in complex scenarios.
@@ V Conclusion
This paper presents MSDet, a novel method for lung nodule detection in CT images, designed to address multi-scale detection challenges. By designed advanced strategies such as TODB, ERD, and PCAM, MSDet enhances detection accuracy, especially for small and occluded nodules. Our model achieves superior performance compared to existing methods, demonstrating a significant improvement in detecting nodules in complex backgrounds. Empirical results on the LUNA16 dataset validate that MSDet outperforms current benchmarks in both precision and recall. This work marks a step forward in medical image analysis, offering valuable insights for future advancements in clinical detection systems.
Supplementary Material
VI Dataset and Evaluation Matrices
In terms of evaluation metrics, the performance of the pulmonary nodule detection model is assessed using several standard measures. Precision quantifies the accuracy of positively predicted instances, aiming to minimize false positives. Recall evaluates the model’s ability to identify all relevant cases, which is crucial for thorough detection. The F1-score merges precision and recall in a single metric, reflecting the balance between detection accuracy and completeness. Average Precision (AP), calculated over the Precision-Recall curve, provides detailed insights into the model’s precision across various recall levels. Mean Average Precision (mAP) averages these AP values across different object classes and sizes, offering a global view of the model’s overall accuracy.
VII Specific preprocessing steps
The lung nodule detection workflow is depicted in Figure 8.
VIII Supplementary Materials – Ablation Study Results
The detailed results of the ablation study are provided in this section. As shown in Supplementary Table II, we present a comprehensive analysis of the model’s performance across different configurations. The study evaluates the impact of various modules, including TODB, ERD, and PCAM, on the model’s precision, recall, and mAP0.5. These results highlight the contribution of each module and demonstrate the significant improvements achieved when all components are integrated. Specifically, the combination of TODB, ERD, and PCAM yields the highest performance, underscoring their importance in enhancing the model’s detection capabilities.
TABLE II: Ablation study of MSDet on the LUNA16 dataset, evaluating the impact of different modules (TODB, ERD, and PCAM) on detection performance.
YOLOv5 TODB ERD PCAM Precision (%) Recall (%) mAP0.5 (%)
83.7 ± 0.09 81.3 ± 0.07 84.7 ± 0.03
88.4 ± 0.04 87.2 ± 0.05 89.1 ± 0.08
88.5 ± 0.02 87.6 ± 0.06 88.9 ± 0.04
87.1 ± 0.09 88.7 ± 0.05 88.3 ± 0.07
91.6 ± 0.08 90.1 ± 0.06 92.8 ± 0.02
91.4 ± 0.04 91.2 ± 0.03 92.5 ± 0.08
91.2 ± 0.09 91.4 ± 0.05 91.9 ± 0.01
97.2 ± 0.06 96.1 ± 0.05 97.3 ± 0.03
TABLE III: Detection performance comparison using different attention modules on the LUNA16 dataset. This table presents precision, recall, F1 score, and mAP results across various IoU thresholds and nodule sizes, demonstrating the superior performance of the PCAM (Ours) module compared to other state-of-the-art attention mechanisms.
Furthermore, Supplementary Table III presents a comparison of different attention mechanisms on the LUNA16 dataset, including SE, CBAM, SA, NAM, CA, ECA, NLNN, GAT, BAM, MAB, CCA, SOCA, and our proposed PCAM module. The table reports precision, recall, F1 score, and mAP results across various IoU thresholds and nodule sizes. The results clearly demonstrate that PCAM outperforms other attention mechanisms in all metrics, achieving the highest precision (93.1%), recall (92.5%), F1 score (92.8%), and mAP0.5 (93.7%) values. Additionally, PCAM consistently shows superior performance across different nodule sizes (mAPS, mAPM, and mAPL), reinforcing its effectiveness in improving multi-scale detection and reducing false positives.
IX Detailed Analysis of Training Progress
The training curves shown in Figure 11 compare the performance of the MSDet model against YOLOv3, YOLOv7, Scaled_YOLOv4, and TPH-YOLOv5 across 200 epochs using metrics such as mAP0.5 and mAP from 0.5 to 0.95. MSDet demonstrates rapid convergence and superior performance, achieving a 60% mAP at 0.5 to 0.95 by epoch 125, highlighting its efficient learning and adaptation capabilities.
The graph presents key metrics including mAP at IoU 0.5, mAP from 0.5 to 0.95, precision, and recall. These metrics illustrate MSDet’s consistently higher precision and recall, maintaining smooth and stable progress. The model’s early high performance, especially in mAP at IoU 0.5, underscores its effective adaptation to complex image data with fewer epochs compared to other models.This enhanced performance is due to MSDet’s sophisticated feature extraction and optimized learning strategies.
Figure 11: Training performance comparison of MSDet with YOLO variants across 200 epochs. The graph demonstrates MSDet’s superior performance with a clear lead in mAP scores from IoU 0.5 to 0.95, underscoring its rapid convergence and higher accuracy in detecting more complex features effectively.